OpenStudy (anonymous):

Solve K = 1/3 (b – 2a) for b ? Please show all work, I'm taking a quiz and really need to understand how to do it :/ sorry - the 1/3 didn't show up the first time i posted

OpenStudy (precal):

quiz is there to assess if you have learned the information. If it is a prequiz, it is to maybe skip instruction, if it is a post quiz it is to see if you understood the instructions

OpenStudy (agent47):

multiply everything by 3 then add 2a to both sides

OpenStudy (anonymous):

k=(1/3)(b-2a) multiply 3 both sides 3k=b-2a add 2a both sides 3k+2a=b-2a+2a 3k+2a=b so, b=3k+2a

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thank you so much!